### v5

@@ 4.1 Group proportions and counterfactual means
Our estimates of the proportions of the two complier groups in the sample use (3.2) in Proposition 5; they are shown in Panel A of Table 5. As expected, they coincide with those in Kline and Walters (2016).
Panel B of Table 5 shows the counterfactual means of test scores for the complier groups, as per Proposition 6. While $\mathbb{E}[Y_{i}(n)|i\in C_{nh}]$ is negative, $\mathbb{E}[Y_{i}(c)|i\in C_{ch}]$ is above 0.10.1 0.10.1 standard deviation—not a negligible value in this field. This suggests that some of the children who enter Head Start would have been at a good preschool otherwise. Kline and Walters (2016) call this pattern the “substitution effect” of Head Start. However, Kline and Walters (2016) do not report estimates of $\mathbb{E}[Y_{i}(n)|i\in C_{nh}]$ and $\mathbb{E}[Y_{i}(c)|i\in C_{ch}]$.
@@
Panel A. Proportions of Response Groups via Proposition 5
@@
Panel B. Counterfactual Means of Test Scores via Proposition 6
$\mathbb{E}[Y_{i}(n)|i\in C_{nh}]$0.027 -0.116 -0.062
@@
Panel C. Treatment Effects via Corollary 2
@@ 4.2 Treatment Effects
To fully measure the substitution effect, one needs to identify $\mathbb{E}\left[Y_{i}(h)|i\in C_{nh}\right]$ and $\mathbb{E}\left[Y_{i}(h)|i\in C_{ch}\right]$. However, we know from Proposition 6 that they are only partially identified by
@@
Kline and Walters (2016) first tried to identify $\mathbb{E}[Y_{i}(h)-Y_{i}(c)|i\in C_{ch}]$ and $\mathbb{E}[Y_{i}(h)-Y_{i}(n)|i\in C_{nh}]$ separately using interactions of the instrument with covariates or experimental sites. They acknowledged the limitations of this approach and resorted to a parametric selection model à la Heckman (1979) instead. They report [^29] See Kline and Walters (2016, Table VIII, column (4), full model). estimates of the local average treatment effects of 0.3700.370 0.3700.370 for $C_{nh}$ and $-0.093$ for $C_{ch}$, with respective standard errors 0.0880.088 0.0880.088 and 0.1540.154 0.1540.154. The resulting point estimate of the difference is quite large, at 0.4630.463 0.4630.463 standard deviation.
Our Corollary 2 provides an alternative approach to separating the two treatment effects. Given that compliers coming from other preschools ($C_{ch}$) had better test scores than compliers not originally in preschools ($C_{nh}$), it seems reasonable to assume that they also have better test scores under Head Start:
@@
This is a “positive selection” that fits within the framework of Corollary 2. It can be derived in a simple model in which preschools, and especially Head Start, improve the outcomes of some students; and students choose schools as a function of their expected outcome. We show in Appendix B that this model generates positive selection under reasonable assumptions. The pooled cohort estimates in Panel C of Table 5 indicate that the upper bound on $\mathbb{E}[Y_{i}(h)-Y_{i}(n)|i\in C_{nh}]$ is 0.280.28 0.280.28 and the lower bound on $\mathbb{E}[Y_{i}(h)-Y_{i}(n)|i\in C_{ch}]$ is 0.090.09 0.090.09. The difference between these two numbers gives an upper bound of 0.190.19 0.190.19 for the difference of these two LATEs, with a standard error of 0.070.07 0.070.07. The testable prediction (3.7) implied by positive selection translates here into the non-negativity of the upper bound; we cannot reject it at any reasonable level. Conversely, negative selection (reverting the inequality (4.1)) would make 0.190.19 0.190.19 a lower bound for the difference of the LATEs. At the same time, it would imply that the lower bound is negative; this is soundly rejected by the data.
Our upper bound of 0.190.19 0.190.19 is much lower than the point estimate reported by Kline and Walters (2016). In fact, our 95% and 99% one-sided confidence intervals for
@@
are $(-\infty,0.308)$ and $(-\infty,0.356)$. We conclude that the 0.4630.463 0.4630.463 estimate in Kline and Walters (2016) may overstate the difference between the two complier groups: it can only be rationalized under negative selection, which is a much less plausible assumption.
@@ 4.3 Expanding Access to Head Start
The children who move from $T_{i}=n$ to $T_{i}=c$ when a slot is vacated by a child who moves to Head Start consitute a $C_{nc}$ group that is ruled out by the $2\times 3$ model. These children increase their grades by $Y_{i}(c)-Y_{i}(n)$, whose average generates a LATE that we denote $\textrm{LATE}_{nc}$. Equation (9) in Kline and Walters (2016, p. 1816) shows that the value of $\textrm{LATE}_{nc}$ is a crucial input in the computation of the MVPF of a Head Start expansion. Identifying it requires either data on offers to all preschools, which Kline and Walters (2016) do not have [^31] See footnote 19 in their paper., or additional modeling assumptions. They used their parametric selection model to construct an estimate for $\textrm{LATE}_{nc}$. Their estimate of $\textrm{LATE}_{nc}=0.294$ results in a high MVPF estimate of 2.022.02 2.022.02 (see Table IX in their paper).
@@
[…]
@@
Now consider the new group of $n\to c$ compliers. It differs from $\tilde{C}_{chc}$ in that its members will not go to a preschool unless they are offered a slot. We show in Appendix B that the structural model that we used in the binary instrument case predicts the following inequality:
@@
Inequalities (4.2) and (4.3) again are “positive selection” assumptions that fall under our Corollary 3.
Since $\tilde{C}_{nn\ast}$ coincides with $C_{nn}$ and $\tilde{C}_{chc}$ is $C_{ch}$, we already know the values of the right-hand sides of both inequalities. Applying the same logic as in Corollary 3 gives us an upper bound for $\textrm{LATE}_{nc}$:
@@
As the MVPF is an increasing function of $\textrm{LATE}_{nc}$, this gives us in turn an upper bound on its value  [^33] Online Appendix F derives the formula for the MVPF in this model.. We obtain $\textrm{LATE}_{nc}\leq 0.164$ and $\textrm{MVPF}\leq 1.55$. These upper bounds are noticeably smaller than the point estimates that result from the parametric selection model of Kline and Walters (2016).
@@
We have shown that the idea of targeting is a useful way to analyze models with multivalued treatments and multivalued instruments. Our paper only analyzed discrete-valued instruments and treatments. Some of the notions we used would extend naturally to continuous instruments and treatments: the definitions of targeting, one-to-one targeting, and positive selection would translate directly. Strict targeting, on the other hand, is less appealing in a context in which continuous values may denote intensities. Our earlier paper (Lee and Salanié, 2018) as well as Mountjoy’s (2022) can be seen as analyzing continuous-instruments/discrete-treatments models. Extending our analysis to models with continuous treatments is an obvious topic for further research. It would also be interesting to apply the partial identification approach of Mogstad, Santos, and Torgovitsky (2018) in our setting. Finally, there has been a surge of recent interest on understanding the properties of OLS and 2SLS estimands when treatment effects vary with the covariates (Blandhol, Bonney, Mogstad, and Torgovitsky, 2022; Słoczyński, 2022; Goldsmith-Pinkham, Hull, and Kolesár, 2022). We believe that the targeting concept and the identifying assumptions explored in this paper should be relevant in this context and that they merit further investigation.
Appendix A Proofs for Section 3
Proof of Proposition 1.
Suppose that for some $t\in\mathcal{T}^{\ast}$, $T_{i}(Z^{\ast}(t))=0$. Then $u_{i0}>\bar{U}(t)+u_{it}$. However, if $z\neq Z^{\ast}(t)$ then $\bar{U}(t)>U_{z}(t)$ under Assumption 4. Therefore $u_{i0}>U_{z}(t)+u_{it}$, and $T_{i}(z)$ cannot be $t$. ∎
Proof of Proposition 2.
Recall that $T_{i}(z)$ maximizes $(U_{z}(t)+u_{it})$ over $t\in\mathcal{T}$. Under strict targeting, $U_{z}(t)$ is $\bar{U}(t)$ if $t\in\mathcal{T}^{\ast}(z)$ and $\underline{U}(t)$ otherwise.
Proof of (i): Since $\bar{U}(t)>\underline{U}(t)$ if $t\in\mathcal{T}^{\ast}$, we have
$V^{\ast}_{i}(z)>\max_{t\in\mathcal{T}^{\ast}(z)}(\underline{U}(t)+u_{it}).$
This implies that
$\displaystyle\max_{t\in\mathcal{T}}(U_{z}(t)+u_{it})$ $\displaystyle=\max\left(\max_{t\in\mathcal{T}^{\ast}(z)}(\bar{U}(t)+u_{it}),\max_{t\not\in\mathcal{T}^{\ast}(z)}(\underline{U}(t)+u_{it})\right)$
$\displaystyle=\max\left(V^{\ast}_{i}(z),\max_{t\in\mathcal{T}}(\underline{U}(t)+u_{it})\right)$
$\displaystyle=\max\left(V^{\ast}_{i}(z),\underline{V}_{i}\right).$
Moreover, if $\underline{V}_{i}=\underline{U}_{\underline{t}_{i}}+u_{i,\underline{t}_{i}}$ is the maximum and $\underline{t}_{i}\in\mathcal{T}^{\ast}(z)$, then a fortiori $\underline{U}_{\underline{t}_{i}}+u_{i,\underline{t}_{i}}>\bar{U}_{\underline{t}_{i}}+u_{i,\underline{t}_{i}}$. This gives a contradiction since $\bar{U}_{t}>\underline{U}_{t}$ for all strictly targeted $t$.
Proof of (ii): If $z\not\in\mathcal{Z}^{\ast}$, then $T^{\ast}(z)$ is empty and $V^{\ast}_{i}(z)=-\infty$. ∎
Proof of Corollary 1.
It follows directly from Proposition 2. ∎
Proof of Proposition 3.
Consider an observation $i$. The set $A_{i}$ of Corollary 1 is a possibly empty subset of $\mathcal{T}^{\ast}$. The top alternative treatment $\underline{t}_{i}$ can be in $A_{i}$ or in $\mathcal{T}\,\setminus\,\mathcal{T}^{\ast}$. If $A_{i}$ has $a$ elements, this allows for $a+\lvert\mathcal{T}\rvert-\lvert\mathcal{T}^{\ast}\rvert=a+q$ values of $\underline{t}_{i}$. Now every pair $(A_{i},\underline{t}_{i})$ fully defines a response-group. Since $\lvert\mathcal{Z}^{\ast}\rvert=\lvert\mathcal{T}^{\ast}\rvert=p$, this gives a total of
$\sum_{a=0}^{p}{p\choose a}(a+q)$
response-groups. Using the identities
$\displaystyle\sum_{a=0}^{b}{b\choose a}$ $\displaystyle=(1+1)^{b}=2^{b}$
$\displaystyle\sum_{a=0}^{b}a{b\choose a}$ $\displaystyle=b\times\sum_{a=0}^{b-1}{b-1\choose a}=b\times 2^{b-1},$
we obtain a total of $(p+2q)\times 2^{p-1}$ types. ∎
Proof of Proposition 4.

### v6

@@ 4.1 Group proportions and counterfactual means
Our estimates of the proportions of the two complier groups in the sample use (3.3) in Proposition˜5; they are shown in Panel A of Table˜6. As expected, they coincide with those in Kline and Walters (2016).
Panel B of Table˜6 shows the counterfactual means of test scores for the complier groups, as per Proposition˜6. While $\mathbb{E}[Y_{i}(n)|i\in C_{nh}]$ is negative, $\mathbb{E}[Y_{i}(c)|i\in C_{ch}]$ is above $0.1$ standard deviation—not a negligible value in this field. This suggests that some of the children who enter Head Start would have been at a good preschool otherwise. Kline and Walters (2016) call this pattern the “substitution effect” of Head Start. However, they do not report estimates of $\mathbb{E}[Y_{i}(n)|i\in C_{nh}]$ and $\mathbb{E}[Y_{i}(c)|i\in C_{ch}]$.
@@
Panel A. Proportions of Response Groups via Proposition˜5
@@
Panel B. Counterfactual Means of Test Scores via Proposition˜6
$\mathbb{E}[Y_{i}(n)|i\in C_{nh}]$ -0.027 -0.116 -0.062
@@
Panel C. Treatment Effects via Proposition˜7
@@ 4.2 Treatment Effects
To fully measure the substitution effect, one needs to identify $\mathbb{E}\left[Y_{i}(h)|i\in C_{nh}\right]$ and $\mathbb{E}\left[Y_{i}(h)|i\in C_{ch}\right]$. However, we know from Proposition˜6 that they are only partially identified by
@@
Kline and Walters (2016) first tried to identify $\mathbb{E}[Y_{i}(h)-Y_{i}(c)|i\in C_{ch}]$ and $\mathbb{E}[Y_{i}(h)-Y_{i}(n)|i\in C_{nh}]$ separately using interactions of the instrument with covariates or experimental sites. They acknowledged the limitations of this approach and resorted to a parametric selection model à la Heckman (1979) instead. They report [^28] See Kline and Walters (2016, Table VIII, column (4), full model). estimates of the local average treatment effects of $0.370$ for $C_{nh}$ and $-0.093$ for $C_{ch}$, with respective standard errors $0.088$ and $0.154$. The resulting point estimate of the difference is quite large, at $0.463$ standard deviation.
Our Proposition˜7 provides an alternative approach to separating the two treatment effects. Given that compliers coming from other preschools ($C_{ch}$) had better test scores than compliers not originally in preschools ($C_{nh}$), it seems reasonable to assume that they also have better test scores under Head Start:
@@
This is a “positive selection” that fits within the framework of Proposition˜7. It can be derived in a simple model in which better students benefit more from Head Start and other preschools; and students choose schools as a function of their expected outcome. We show in Appendix C.1 that this model generates positive selection under reasonable assumptions. The pooled cohort estimates in Panel C of Table˜6 indicate that the upper bound on $\mathbb{E}[Y_{i}(h)-Y_{i}(n)\mid i\in C_{nh}]$ is $0.28$, while the lower bound on $\mathbb{E}[Y_{i}(h)-Y_{i}(c)\mid i\in C_{ch}]$ is $0.09$. [^29] Under monotone treatment response, $Y_{i}(h)\geq Y_{i}(n)$, the lower bound on $\mathbb{E}[Y_{i}(h)-Y_{i}(n)\mid i\in C_{nh}]$ would be zero, yielding a 95% confidence interval $0\leq\mathbb{E}[Y_{i}(h)-Y_{i}(n)\mid i\in C_{nh}]\leq 0.36$. The difference between these two numbers gives an upper bound of $0.19$ for the difference of these two LATEs, with a standard error of $0.07$. Negative selection (reverting the inequality (4.1)) would make $0.19$ a lower bound for the difference of the LATEs. At the same time, it would imply that the lower bound is negative; this is soundly rejected by the data.
Our upper bound of $0.19$ is much lower than the point estimate reported by Kline and Walters (2016). In fact, our 95% and 99% one-sided confidence intervals for
@@
are $(-\infty,0.308)$ and $(-\infty,0.356)$. We conclude that the $0.463$ estimate in Kline and Walters (2016) may overstate the difference between the two complier groups: it can only be rationalized under negative selection, which is a much less plausible assumption.
@@ 4.3 Expanding Access to Head Start
The children who move from $T_{i}=n$ to $T_{i}=c$ when a slot is vacated by a child who moves to Head Start consitute a $C_{nc}$ group that is ruled out by the $2\times 3$ model. These children increase their grades by $Y_{i}(c)-Y_{i}(n)$, whose average generates a LATE that we denote $\textrm{LATE}_{nc}$. Equation (9) in Kline and Walters (2016, p. 1816) shows that the value of $\textrm{LATE}_{nc}$ is a crucial input in the computation of the MVPF of a Head Start expansion. Identifying it requires either data on offers to all preschools, which Kline and Walters (2016) do not have [^31] See footnote 19 in their paper., or additional modeling assumptions. They used their parametric selection model to construct an estimate for $\textrm{LATE}_{nc}$. Their estimate of $\textrm{LATE}_{nc}=0.294$ results in a high MVPF estimate of $2.02$ (see Table IX in their paper).
@@
$u_{ih}-u_{in}$$u_{ic}-u_{in}$$\tilde{C}_{nhc}$$\tilde{C}_{nhn}$$\tilde{C}_{nnc}$$\tilde{C}_{nnn}$$\tilde{C}_{hhh}$$\tilde{C}_{ccc}$$\tilde{C}_{hhc}$$\tilde{C}_{chc}$$P_{n}$$P_{h}$$P_{c}$
@@
Now consider the new group of $n\to c$ compliers. It differs from $\tilde{C}_{chc}$ in that its members will not go to a preschool unless they are offered a slot. We show in Appendix C.2 that the structural model that we used in the binary instrument case predicts the following inequality:
@@
Inequalities (4.2) and (4.3) again are “positive selection” assumptions that fall under our Proposition˜10.
Since $\tilde{C}_{nn\ast}$ coincides with $C_{nn}$ and $\tilde{C}_{chc}$ is $C_{ch}$, we already know the values of the right-hand sides of both inequalities. Applying the same logic as in Proposition˜10 gives us an upper bound for $\textrm{LATE}_{nc}$:
@@
As the MVPF is an increasing function of $\textrm{LATE}_{nc}$, this gives us in turn an upper bound on its value  [^33] Online Appendix G derives the formula for the MVPF in this model.. We obtain $\textrm{LATE}_{nc}\leq 0.164$ and $\textrm{MVPF}\leq 1.55$. These upper bounds are noticeably smaller than the point estimates that result from the parametric selection model of Kline and Walters (2016).
@@
We have shown that the idea of targeting is a useful way to analyze models with multivalued treatments and multivalued instruments. The testable implications that we derived suggest a natural three-step procedure to elucidate targeting patterns in the data, supplementing any qualitative information provided by the empirical context.
Step 1: Screening Candidates.
For each instrument-target pair $(z,t)\in\mathcal{Z}\times\mathcal{T}$, test the hypothesis $H_{0,zt}:S_{zt}\leq 0$ against $H_{1,zt}:S_{zt}>0$, where
$S_{zt}=P(0\mid z)+\max_{z^{\prime\prime}\in\mathcal{Z},z^{\prime\prime}\neq z}P(t\mid z^{\prime\prime})-1.$
Let $\widehat{\mathcal{V}}$ denote the set of surviving pairs:
$\widehat{\mathcal{V}}\equiv\{(z,t)\in\mathcal{Z}\times\mathcal{T}:H_{0,zt}\ \text{is not rejected}\}.$
If for a fixed $z$, no pair $(z,t)$ survives in $\widehat{\mathcal{V}}$, we reject the hypothesis that $z$ targets any treatment in $\mathcal{T}$.
Step 2: Checking Compatibility
For every pair of candidates $(z,t)$ and $(z^{\prime},t^{\prime})$ in $\widehat{\mathcal{V}}$ with $t\neq t^{\prime}$, we test $H_{0,c}:S_{c}\leq 0$, where
$S_{c}=P(t^{\prime}\mid z)+P(t\mid z^{\prime})-1.$
A rejection implies that $(z,t)$ and $(z^{\prime},t^{\prime})$ cannot hold simultaneously.
Step 3: Testing for Universal Targeting
The set of valid mappings that survive Steps 1 and 2 may or may not be consistent with universal targeting. One can further use the inequalities in (2.6) to formally test it.
As usual, one should be careful of controlling the size of this three-step procedure by applying standard multiple-testing corrections within each step and/or bootstrapping. We leave this for further research.
@@
Our paper only analyzed discrete-valued instruments and treatments. Some of the notions we used would extend naturally to continuous instruments and treatments: the definitions of targeting, one-to-one targeting, and positive selection would translate directly. Strict targeting, on the other hand, is less appealing in a context in which continuous values may denote intensities. Our earlier paper (Lee and Salanié, 2018) as well as Mountjoy’s (2022) can be seen as analyzing continuous-instruments/discrete-treatments models. Extending our analysis to models with continuous treatments is an obvious topic for further research. It would also be interesting to apply the partial identification approach of Mogstad, Santos, and Torgovitsky (2018) in our setting. Finally, there has been a surge of recent interest on understanding the properties of OLS and 2SLS estimands when treatment effects vary with the covariates (Blandhol, Bonney, Mogstad, and Torgovitsky, 2022; Słoczyński, 2022; Goldsmith-Pinkham, Hull, and Kolesár, 2024). We believe that the targeting concept and the identifying assumptions explored in this paper should be relevant in this context and that they merit further investigation.
Appendix A Proofs for Section˜2
Proof of Proposition˜2.
As is explained in the main text, it is a direct consequence of Proposition˜1 under one-to-one targeting. ∎
Proof of Proposition˜3.
(i): Suppose that $z$ is not a targeting instrument. Then $T_{i}(z)=\underline{t}_{i}$ (which does not depend on $z$ under strict targeting). If $z$ is targeting but $z\not\in A_{i}$, then $T_{i}(z)\neq z$; since $z$ targets $z$ uniquely, $t^{\ast}_{i}(z)=z$ and $T_{i}(z)$ can only be $\underline{t}_{i}$.
(ii): If $\underline{t}_{i}$ is in $\mathcal{T}^{\ast}$, we must have $\underline{V}_{i}=\underline{U}(\underline{t}_{i})+u_{i\underline{t}_{i}}<\bar{U}(\underline{t}_{i})+u_{i\underline{t}_{i}}=V^{\ast}_{i}(\underline{t}_{i})$; therefore $T_{i}(\underline{t}_{i})=\underline{t}_{i}$ and $\underline{t}_{i}$ is in $A_{i}$.
(iii): We already know that each observation $i$ is in a $C(A,\tau)$ set. Then $T_{i}(t)=t$ over $A$ and $T_{i}(z)=\tau$ for all other $z$. This defines an elemental response-group. If $i$ also belongs to $C(A^{\prime},\tau^{\prime})$, then $A$ must equal $A^{\prime}$ since they are both defined as the subset of $\mathcal{T}^{\ast}$ where $T_{i}(t)=t$. Moreover, $T_{i}(0)=\tau=\tau^{\prime}$ since $0\not\in\mathcal{T}^{\ast}$. Therefore $\tau=\tau^{\prime}$ and the various $C(A,t)$ form a partition of the set of observations. ∎
Appendix B Proofs for Section˜3
Proof of Lemma˜1.
We start from the definition:
$\displaystyle\bar{E}_{z}(t)$ $\displaystyle=\mathbb{E}(Y_{i}{\rm 1\kern-3.99994pt1}(T_{i}=t)\mid Z_{i}=z)$
$\displaystyle=\sum_{C}\mathbb{E}(Y_{i}{\rm 1\kern-3.99994pt1}(T_{i}=t)\mid i\in C,Z_{i}=z)\Pr(i\in C\mid Z_{i}=z)$
$\displaystyle=\sum_{C}\mathbb{E}(Y_{i}(t)\mid T_{i}=t,i\in C,Z_{i}=z)\Pr(T_{i}=t,i\in C\mid Z_{i}=z)$
$\displaystyle=\sum_{C}\mathbb{E}(Y_{i}(t)\mid T_{i}(z)=t,i\in C,Z_{i}=z)\Pr(T_{i}(z)=t,i\in C\mid Z_{i}=z),$
where the sum runs over all complier groups and the third equality uses Assumption 5(i).
Now if $C=C_{R}$ and $R(z)\neq t$, $\Pr(T_{i}(z)=t,i\in C\mid Z_{i}=z)$ is zero; and if $R(z)=t$, then we can drop the $T_{i}(z)=t$” terms:
$\mathbb{E}(Y_{i}(t)\mid T_{i}(z)=t,i\in C,Z_{i}=z)=\mathbb{E}(Y_{i}(t)\mid i\in C,Z_{i}=z),$
which equals $\mathbb{E}(Y_{i}(t)\mid i\in C)$ by Assumption 5(ii); and
$\Pr(T_{i}(z)=t,i\in C\mid Z_{i}=z)=\Pr(i\in C\mid Z_{i}=z),$
which equals $\Pr(i\in C)$ by Assumption 5(ii). Therefore we can rewrite the sum as
$\bar{E}_{z}(t)=\sum_{C_{R}\;|\;R(z)=t}\mathbb{E}(Y_{i}(t)\;|i\in C_{R})\Pr(i\in C_{R}).$
This proves the first result in the lemma.
The second result in the lemma follows the same logic by observing that $\Pr(T_{i}=t\mid Z_{i}=z)=\mathbb{E}[{\rm 1\kern-3.99994pt1}(T_{i}=t)\mid Z_{i}=z]$, which is equivalent to the first result where $Y_{i}=1$ (and thus $Y_{i}(t)=1$). ∎
Proof of Proposition˜4.